OpenStudy (volleyballlover55):

Part A: Find a rational number that is between 9.5 and 9.7. Explain why it is rational. Part B: Find an irrational number that is between 9.5 and 9.7. Explain why it is irrational. Include the decimal approximation of the irrational number to the nearest hundredth.

OpenStudy (volleyballlover55):

@jim_thompson5910

OpenStudy (zarkon):

if a and b are rational numbers then look at (a+b)/2

jimthompson5910 (jim_thompson5910):

Part B) To find an irrational number between x and y, find the geometric mean M \[\Large M = \sqrt{x*y}\] M is irrational if and only if x or y are not perfect squares

OpenStudy (zarkon):

I think you mean M is irrational if and only if x*y is not perfect square

jimthompson5910 (jim_thompson5910):

No I meant it like this \[\Large M = \sqrt{x*y}\] \[\Large M = \sqrt{x}*\sqrt{y}\] In the second equation, sqrt(x) is only rational if and only if x is a perfect square. Same for y.

jimthompson5910 (jim_thompson5910):

so if x wasn't a perfect square, or y wasn't, then M would be irrational

OpenStudy (zarkon):

x=12, y=75 x*75=900=30^2

jimthompson5910 (jim_thompson5910):

oh true. I didn't think of that. So yeah x*y not being a perfect square is a better condition

OpenStudy (zarkon):

you could also use a known irrational...\(\pi= 3.141592...\) then \[\pi/100=.03141592...\] then \[9.5+\pi/100=9.53141592...\] which is irrational (it has a non-terminating non-repeating decimal expansion)